In an back yard vineyard in Napa valley with 10 grape vines (just right), rain in the spring and dry through summer, roughly Please use BAYES rule! The other examples aren't showing this problem sovled using Bayes rule.

solve using Bayes rule only! no z-scores allowed.
Answer #1

Solution:

Given that, The yield of for each vine is distributed goughly binomial with N = 900 P = 0.7 In a drought the yield is binomia. We know that for a right year X Sem Binomial (N=900, P=0.7) using Normal approximation to Binomial 1 x follows Normal with$ 100 when ( a bottle is sold foot weatheor is dory We have to find, P(X7631 and dory weather) NOW, P(X7631 and dory weatheonP (x>631 and day weathe I and day weatheon) = P(x>631 | dory weathear ) * P(dory weather) = 0.0057 *0.5 = 0.0029 ,: The porothen P ( x x6 31 and wer weather) - P (x>631 wet weather )*e (wer weather) = 0.0000 *0.2 = 0.0000 . The probability that wine